The width of a rectangle is 3cm less than its length. The rectangle has area of 40cm^2. What is the length of the rectangle?
Please help Im not very good at word problems!!

Answer by stanbon(75887)

The width of a rectangle is 3cm less than its length. The rectangle has area of 40cm^2. What is the length of the rectangle?
------
Let length be "x":
Then width is "x-3":
---
Area = length * width
40 = x(x-3)
x^2-3x-40 = 0
(x-8)(x+5) = 0
Positive solution:
x = 8 cm (length)
x-3 = 5 cm (width)
------------
Cheers,
Stan H.
=============